A bullet is fired perpendicular upwards into the air and leaves the barrel at 200m/s. answer the following questions:

1.How long would it take before bullet has an upward velocity at 80 m/s?
2.How long would it take before the bullet reaches its maximum height?
3.What will the velocity of the bullet be after 15 s?
Chat started
1
Expert's answer
2021-03-16T11:35:48-0400

1)

"v=v_0+gt,""t=\dfrac{80\ \dfrac{m}{s}-200\ \dfrac{m}{s}}{-9.8\ \dfrac{m}{s^2}}=12.24\ s."

2)

"0=v_0+gt,""t=\dfrac{-v_0}{g}=\dfrac{-200\ \dfrac{m}{s}}{-9.8\ \dfrac{m}{s^2}}=20.41\ s."

3)

"v(t=15\ s)=200\ \dfrac{m}{s}+(-9.8\ \dfrac{m}{s^2})\cdot15\ s=53\ \dfrac{m}{s}."